OpenStudy (anonymous):

Find the derivative of f(x) = -12x2 + 9x at x = 6 Please show me how to do this!

OpenStudy (anonymous):

f(x)=-24+9(6)

OpenStudy (anonymous):

and then 9(6)=54 and so it should look like f(x)=30

OpenStudy (john_es):

\[f'(x)=-12\cdot2\cdot x^{2-1}+9=-24x+9\]Then evaluate in x=6.

OpenStudy (anonymous):

My answers are -112.5 -135 -90 -108 Thats why I'm confused

OpenStudy (john_es):

You substitue x=6, \[-144+9=-135\]